If two (2) molecules of 2-phosphoglycerate is oxidized via glycolysis, the PDC, and the CAC in a kidney cell, how many net molecules of ATP (plus GTP) will be produced by oxidative phosphorylation and substrate level phosphorylation? Assume that this particular ATP synthase has 9 c-subunits and that the phosphate translocate symporter is utilized. (Net ATP = total ATP produced - total ATP used up.)
Answer #1

When one molecule of 3-phosphoglycerate is converted to one molecule of pyruvate 1 ATP is formed by substrate level phosphorylation. When pyruvate is oxidized to acetyl CoA one NADH is produced. When one acetyl CoA enters into TCA, 3 NADH, 1 FADH2 and 1 ATP (again substrate level phosphorylation) is produced.

So total 4 NADH, 1 FADH2 and 2 ATP are produced.

Consider NADH = 2.5 ATP and FADH2 = 1.5 ATP

Net ATP produced including substrate level phosphorylation = 4×2.5 + 1.5 + 2 = 13.5 ATP

Out of 13 ATP, two molecules of ATP are produced by substrate level phosphorylation.

So ATP produced only by oxidative phosphorylation = 13.5 - 2 = 11.5